# Multilayered Intelligent Reflecting Surface for Long-Range Underwater Acoustic Communication

### v1

@@ Abstract
This article introduces a multilayered acoustic reconfigurable intelligent surface (ML-ARIS) architecture designed for the next generation of underwater communications. ML-ARIS incorporates multiple layers of piezoelectric material in each acoustic reflector, with the load impedance of each layer independently adjustable via a control circuit. This design increases the flexibility in generating reflected signals with desired amplitudes and orthogonal phases, enabling passive in-phase and quadrature (IQ) modulation using a single acoustic reflector. Such a feature enables precise beam steering, enhancing sound levels in targeted directions while minimizing interference in surrounding environments. Extensive simulations and tank experiments were conducted to verify the feasibility of ML-ARIS. The experimental results indicate that implementing IQ modulation with a multilayer structure is indeed practical in real-world scenarios, making it possible to use a single reflection unit to generate reflected waves with high-resolution amplitudes and phases.
Index Terms:
Acoustic reconfigurable intelligent surface (ARIS), multilayer structure, in-phase and quadrature (IQ) modulation, underwater acoustic communication.
@@ I Introduction
The passive operation of RIS offers superior power conservation compared to active communication systems, making it highly sustainable for extreme environments, such as underwater settings where battery replacement is impractical. However, existing RIS research predominantly focuses on radio frequency (RF) communications, which face significant range limitations in aquatic settings due to substantial absorption losses in water [4].
As a result, acoustic communication has been the principal method for mid-range and long-range underwater communication. Due to the fundamental differences between RF and acoustic signals, current RF-RIS designs are unsuitable for direct application in underwater acoustic RIS (UA-RIS). Currently, only a few studies explored the design of UA-RIS [5, 6], highlighting a significant opportunity for advancements and breakthroughs in the field.
Our field experiments have confirmed the potential of UA-RIS to enhance the range and data rate of underwater acoustic communications [7]. Nevertheless, transitioning UA-RIS from theory concepts to practical application still needs to overcome a series of unique challenges: {adjustwidth}-0.71cm0cm
@@
To address these challenges, we introduce a multilayered acoustic reconfigurable intelligent surface (ML-ARIS) featuring a stacked architecture to enable flexible acoustic reflection. In ML-ARIS, the reflector comprises multiple stacked PZT disks, each with independently adjustable load impedance via a control circuit. This design increases the flexibility in generating reflected signals with desired amplitudes and orthogonal phases. By programming the control circuit, the superposed wave from each reflector can have any desired phase and amplitude. Consequently, ML-ARIS can execute passive in-phase and quadrature (IQ) modulation on a single acoustic reflector, enabling precise beam steering and significantly boosting wave intensity toward targeted areas.
@@
Unlike multi-bit coding RF-RIS, which typically produces a very limited number of phase shift states and lacks control over the strength of the reflected signal [8, 9], ML-ARIS with IQ modulation can generate a high-resolution reflection coefficient. Consequently, ML-ARIS is capable of performing advanced array processing methods such as robust capon beamforming (RCB) and minimum variance distortionless response (MVDR). These techniques enhance the intensity of reflected waves in targeted directions while mitigating unintended interference in surrounding environments.
@@
In addition to the multilayer structure, an enhanced matching network is implemented to address the varying impedance of piezoelectric materials in dynamic environments. This network features a three-tier subnetwork architecture, with each tier comprising a high-pass L-type matching circuit. The interconnection between the tiers is managed by a microcontroller (MCU), allowing adaptation to changes in the reflector’s impedance caused by temperature fluctuations and the incident angle of acoustic waves. To ensure optimal performance, the tiers operate in a cascaded manner rather than independently. The parameters of all components in the matching circuit, including capacitance and inductance, are determined by solving an optimization problem.
@@
We conducted extensive experiments in a tank environment, as well as simulations using the COMSOL platform, to evaluate the performance of ML-ARIS around 28 kHz and 41 kHz. The tests demonstrated that with the proposed matching circuit, the magnitude of the reflection coefficient at the target load remains below 0.15 as the water temperature ranges from 9°C to 22°C and the acoustic signal frequency varies between 27.5 kHz and 28.5 kHz at different water depths. Both the simulation and tank experiment results confirm that the stacked architecture with IQ modulation can effectively manipulate the amplitude and phase of reflected waves on a single reflection unit. Compared to multi-bit coding schemes, ML-ARIS with IQ modulation generates significantly lower side lobes in the reflected beam, enabling more devices to share the acoustic channel without causing interference issues. In particular, our work makes three key contributions: {adjustwidth}-0.71cm0cm
@@
A multilayer structure-based ARIS is proposed for underwater acoustic communications. ML-AIRS enables IQ modulation on each reflector, allowing for creation of reflected signal with tailored amplitude and phase, enabling precise beam steering through advanced array processing.
@@
We introduce a specialized matching network to accommodate variations in the reflector’s impedance caused by environmental changes. This ensures consistent impedance matching along the transmission line, crucial for accurate IQ modulation.
@@ II Related Work
In recent years, extensive research has been undertaken to optimize RF-RIS performance, address implementation challenges, and explore its applicability across diverse communication scenarios. For instance, in [10], three distinct free-space path loss models were developed. These models analyze the relationship between the path loss in RIS-assisted wireless communications and variables such as the distances to the RIS, the RIS size, and the radiation patterns of reflected waves. Their models are validated through experiments with three fabricated RIS, which paves the way for future theoretical insights and practical innovations in the field of RF-RIS.
@@
The work presented in [11] investigates the energy efficiency of RIS-assisted communication systems. By simultaneously optimizing the transmitter power and the phase shifts of RIS elements, their theoretical analysis suggests that an RIS-assisted network could achieve up to 300% higher energy efficiency compared to conventional multi-antenna amplify-and-forward relaying systems, without compromising quality of service. Additionally, [12] discusses deployment strategies for RIS in millimeter-wave (mmWave) multi-cell networks. The study introduces environment-aware algorithms to optimize RIS placement, enhancing coverage and capacity while mitigating the signal blockages typical in mmWave communications.
@@
In [13] a survey of RF-RIS hardware development is provided. The survey highlights the challenges and innovations in designing cost-effective RIS elements. The role of metametamaterials and programmable metasurfaces in achieving dynamic control over EM waves is emphasized. Specific techniques for achieving flexible phase shifts using varactor diodes are examined in [14], [15], and [16]. These works involve placing a varactor diode at the radiating edge of a patch antenna and adjusting its reverse voltage to electronically control the RIS’s phase response, allowing for a broad tunable phase control while maintaining a high reflection efficiency.
@@
Currently, only a limited number of studies have investigated the application of RIS techniques in underwater environments. For instance, the research presented in [17] demonstrates the feasibility of using backscattered signals for long-distance passive communications. In this work, a Van Atta array is constructed to automatically align the direction of reflected acoustic waves with the incident waves. By switching the load of each reflector between short and open circuit states, the polarization of the reflected wave is toggled. Consequently, the strength of the received signal, which is the superposition of emitted and reflected waves, exhibits variations. Sea experiments indicate that the Van Atta backscatter array can achieve a round-trip communication range exceeding 300 meters with a bit error rate (BER) of $10^{-3}$.
@@
In [5], the enhancement of communication performance using UA-RIS is examined. The design for each reflection unit incorporates a piezoelectric coil positioned between two metal plates, with the resonant frequency adjusted by the reverse voltage applied to a varactor. Meanwhile, [6] explores the design of UA-RIS capable of responding to a wide range of frequencies. This work addresses two specific types of dispersion: element dispersion within each piezoelectric reflector and array dispersion throughout the acoustic RIS. By counteracting these dispersions, the directivity of the reflected waves becomes frequency-independent, allowing wideband signals to achieve uniform beam patterns across various frequency components. The efficacy of this new architecture is validated through Bellhop and COMSOL simulators.
@@ III Motivation of Multilayered UA-RIS
This section presents the motivation for our work. We begin by discussing the importance of simultaneously manipulating both the amplitude and phase of reflected waves in UA-RIS applications. Afterward, the inherent challenges in achieving flexible reflection control are clarified. Finally, we briefly review our previous solution for implementing such control and highlight its limitations.
@@ III-A Necessity of Flexible Reflection Control
The underwater acoustic channel is a precious resource. It is essential not only for manmade facilities but also for marine animals that rely on acoustic signals for hunting, communication, and navigation [18]. However, the absorption attenuation of sound signals increases significantly with frequency, limiting the available bandwidth for mid-range and long-range acoustic applications to just tens of kilohertz. As the number of marine facilities increases, efficiently utilizing the underwater channel without causing harm to surrounding environments presents a critical challenge for future acoustic systems.
To meet the above objective, it is crucial to design RIS-assisted underwater acoustic communication systems that not only optimize signal strength at the main lobe but also minimize interference from side lobes. The typical 1-bit and 2-bit coding schemes used in RF-RIS may fall short, as the phase resolution of RIS elements is a significant determinant of system performance. According to array processing principles [19], an RIS with arbitrary reflection coefficients can increase the directional gain of the reflected wave by approximately 23.46% ($\sim$ 0.9 dB) over a 2-bit coding RIS and diminish side lobe interference by around 50% ($\sim$ 3 dB).
@@
III-B Challenges of Flexible Reflection Control
In RF-RIS, each unit can flexibly control the phase of the reflection coefficient by adjusting the load reactance. This is accomplished by integrating varactor diodes into the load and modifying the reverse bias voltage applied to the varactor’s p-n junction [16, 20]. However, this approach is impractical for UA-RIS due to the low frequencies associated with acoustic signals. Specifically, the reflection coefficient, denoted as $\Gamma$, in a transmission line is defined as:
@@ III-B Challenges of Flexible Reflection Control
Based on the above equation, if a reflector is designed to adjust the phase of  $\Gamma$ from --45° to --90°, the load impedance should change from $-j$121 $\Omega$ to $-j$50 $\Omega$, where $j$ represents the imaginary part of the impedance contributed by the reactive components. In this scenario, if an RF-RIS operates at the 2.4 GHz frequency band, the capacitance of the load capacitor needs to vary from 0.55 pF to 1.33 pF. This range falls within the capability of many commercial varactor diodes [21, 22]. In contrast, if the frequency is reduced to 25 kHz for long-distance acoustic communication, the required capacitance of the load capacitor must change from 52 nF to 128 nF to achieve the same phase adjustment. This range significantly exceeds the capabilities of most varactor diodes, which typically have capacitance values below 1 nF.
III-C Previous Solution of Flexible Reflection Control
@@
In our previous work [7], we introduced an IQ modulation-based RIS for synthesizing acoustic waves with customizable amplitude and phase. That approach required generating in-phase and quadrature components using two distinct reflectors, which significantly reduced performance relative to the multilayer structure proposed in this paper. In the following, we provide a detailed comparison of the two methods.
Figure 1: IQ modulation implemented with separated reflectors.
@@
Fig. 1 illustrates the IQ modulation scheme from [7]. In this configuration, two adjacent reflectors sharing the same Y-coordinate are paired. For example, reflector 5 is paired with reflector 6. When paired, these reflectors collectively act as a single reflector on reflection plane 1. This is possible because their projected positions along the Y-axis are aligned, ensuring that positional differences do not introduce additional phase shifts on that plane. During operation, the paired reflectors produce orthogonal reflection coefficients, allowing the synthesis of reflected waves with arbitrary phase waves on plane 1 through appropriate adjustments of the amplitude ratio between these orthogonal components.
@@
Using the above approach, two groups of reflectors indicated by purple and blue in Fig. 1 are formed. Each group consists of four reflector pairs, forming four collective reflectors along the Y-axis. These collective reflectors can generate any desired amplitude and phase of the reflection coefficients. By varying the phase differences among neighboring collective reflectors, the angle of reflection plane 1 can be effectively altered.
III-D Disadvantages of Previous Work
@@
In an ideal scenario, the reflected wave would be superposed along a single line oriented toward the target, rather than spread over a plane. As illustrated in Fig. 1, this is achieved by configuring two sets of linear arrays along the X and Y axis to form two reflection planes. The intersection of these planes defines the directivity of this two-dimensional array.
@@
However, for the UA-RIS proposed in our previous research [7], all reflection coefficients are determined solely by the angle of the reflection plane 1 and the incident wave, leaving no additional DoF to establish a second reflection plane. Implementing such an additional plane would require pairing reflectors along the X-axis and Y-axis simultaneously, which is not feasible under that design. As a result, performing IQ modulation with separate reflectors can only achieve one-dimensional gain, offering no directivity improvement in the orthogonal dimension. For an $N\!\times\!N$ array, this reduces the performance to that of $N/2$ independent linear arrays, each with $N$ acoustic reflectors.
@@
To improve the performance of UA-RIS, a new structure is needed to flexibly control the phase and the amplitude of reflected waves. The criteria for this structure are: (a) The design should achieve optimal or near-optimal main lobe gain and side lobe suppression, approaching the performance of an ideal two-dimensional RIS. (b) It should readily adapt to varying environmental conditions, such as variation in intensity and angle of the incident wave, temperature changes, and deployment depths, thereby expanding the application range of UA-RIS. (c) The new structure should be economically feasible, enabling large-scale deployment.

### v2

# ML-ARIS: Multilayer Underwater Acoustic Reconfigurable Intelligent Surface with High-Resolution Reflection Control

@@ Abstract
This article introduces a multilayered acoustic reconfigurable intelligent surface (ML-ARIS) architecture designed for the next generation of underwater communications. ML-ARIS incorporates multiple layers of piezoelectric material in each acoustic reflector, with the load impedance of each layer independently adjustable via a control circuit. This design increases the flexibility in generating reflected signals with desired amplitudes and orthogonal phases, enabling passive synthetic reflection using a single acoustic reflector. Such a feature enables precise beam steering, enhancing sound levels in targeted directions while minimizing interference in surrounding environments. Extensive simulations and tank experiments were conducted to verify the feasibility of ML-ARIS. The experimental results indicate that implementing synthetic reflection with a multilayer structure is indeed practical in real-world scenarios, making it possible to use a single reflection unit to generate reflected waves with high-resolution amplitudes and phases.
@@ I Introduction
Compared with active acoustic relays or repeaters, passive operation of RIS offer a fundamentally different operational trade-off: they shape the acoustic propagation environment without generating new acoustic signals. This makes RIS particularly attractive in scenarios where power availability, acoustic self-noise, or system visibility are critical constraints. Representative applications include long-duration seabed sensor networks and observatories, where battery replacement is costly [22]; infrastructure-assisted coverage extension for autonomous underwater vehicle (AUV) operations, where passive surfaces can redirect energy into desired regions; and covert or environmentally sensitive deployments, where minimizing acoustic emissions is desirable [14]. In these settings, RIS complements active systems by providing a low-power, low-complexity means of enhancing link reliability and spatial selectivity, rather than replacing transmitters or receivers. This motivates the need for advanced reflection control and environmental adaptability.
Existing RIS research predominantly focuses on radio frequency (RF) communications, which face significant range limitations in aquatic settings due to substantial absorption losses in water [20]. As a result, acoustic communication has been the principal method for mid-range and long-range underwater communication. Due to the fundamental differences between RF and acoustic signals, current RF-RIS designs are unsuitable for direct application in underwater acoustic RIS (UA-RIS). Currently, only a few studies explored the design of UA-RIS [21, 27, 12], highlighting a significant opportunity for advancements and breakthroughs in the field.
Our field experiments have confirmed the potential of UA-RIS to enhance the range and data rate of underwater acoustic communications [13]. Nevertheless, transitioning UA-RIS from theory concepts to practical application still needs to overcome a series of unique challenges:
@@
To address these challenges, we introduce a multilayered acoustic reconfigurable intelligent surface (ML-ARIS) featuring a stacked architecture to enable flexible acoustic reflection. In ML-ARIS, the reflector comprises multiple stacked PZT disks, each with independently adjustable load impedance via a control circuit. This design increases the flexibility in generating reflected signals with desired amplitudes and orthogonal phases. By programming the control circuit, the superposed wave from each reflector can achieve a wide and continuously tunable range of phase and amplitude values, subject to the physical and circuit constraints of the multilayer reflector. Consequently, ML-ARIS can execute passive synthetic reflection on a single acoustic reflector, enabling precise beam steering and significantly boosting wave intensity toward targeted areas.
@@
Unlike multi-bit coding RF-RIS, which typically produces a very limited number of phase shift states and lacks control over the strength of the reflected signal [11, 32], ML-ARIS with synthetic reflection can generate a high-resolution reflection coefficient. Consequently, ML-ARIS is capable of performing advanced array processing methods such as robust capon beamforming (RCB) and minimum variance distortionless response (MVDR). These techniques enhance the intensity of reflected waves in targeted directions while mitigating unintended interference in surrounding environments.
@@
In addition to the multilayer structure, an enhanced matching network is implemented to address the varying impedance of piezoelectric materials in dynamic environments. It utilizes a three-tier subnetwork architecture, each tier comprising a high-pass L-type matching circuit. A microcontroller (MCU) coordinates the interconnections between tiers, adapting to changes in the reflector’s impedance due to temperature fluctuations and the incident angle of acoustic waves. For optimal performance, the tiers operate in a cascaded manner rather than independently. The parameters of all components in the matching circuit, including capacitance and inductance, are determined by solving an optimization problem.
@@
We conducted extensive experiments in a tank environment, as well as simulations using the COMSOL platform, to evaluate the performance of ML-ARIS at around 28 kHz and 41 kHz. The tests demonstrated that with the proposed matching circuit, the magnitude of the reflection coefficient at the target load remains below 0.15 as the water temperature ranges from 9C to 22C and the acoustic signal frequency varies between 27.5 kHz and 28.5 kHz at different water depths. Both the simulation and tank experiment results confirm that the stacked architecture with synthetic reflection effectively manipulates the amplitude and phase of reflected waves on a single reflection unit. Compared to multi-bit coding schemes, ML-ARIS with synthetic reflection generates significantly lower side lobes in the reflected beam, enabling more devices to share the acoustic channel without causing interference. In particular, our work makes three key contributions:
@@
A multilayer structure-based ARIS is proposed for underwater acoustic communications. ML-ARIS enables synthetic reflection on each reflector, allowing for creation of reflected signal with tailored amplitude and phase, enabling precise beam steering through advanced array processing.
@@
We introduce a specialized matching network to accommodate variations in the reflector’s impedance caused by environmental changes. This ensures consistent impedance matching along the transmission line, crucial for accurate synthetic reflection.
@@ II Related Work
In recent years, extensive research has been undertaken to optimize RF-RIS performance, address implementation challenges, and explore its applicability across diverse communication scenarios. For instance, in [24], three distinct free-space path loss models were developed to analyze the relationship between the path loss in RIS-assisted wireless communications and variables such as the distances, RIS size, and radiation patterns of reflected waves. These models were validated through experiments with fabricated RIS, paving the way for future theoretical insights and practical innovations in the field.
@@
The work presented in [9] investigated the energy efficiency of RIS-assisted communication systems. By optimizing both the transmitter power and the phase shifts of RIS elements, their theoretical analysis suggests that an RIS-assisted network could achieve up to 300% higher energy efficiency compared to conventional multi-antenna amplify-and-forward relaying systems, without compromising quality of service. Additionally, [19] discussed deployment strategies for RIS in millimeter-wave (mmWave) multi-cell networks. This study introduced environment-aware algorithms to optimize RIS placement, enhancing coverage and capacity while mitigating the signal blockages typical in mmWave communications.
@@
In [10], a survey of RF-RIS hardware development is provided, highlighting challenges and innovations in designing cost-effective RIS elements. The role of metametamaterials and programmable metasurfaces in achieving dynamic control over EM waves was emphasized. Techniques for achieving flexible phase shifts using varactor diodes were examined in [5], [23], and [18]. These works involve placing a varactor diode at the radiating edge of a patch antenna and adjusting its reverse voltage to electronically control the RIS’s phase response, allowing for a broad tunable phase control.
@@
Currently, only a limited number of studies have investigated the application of RIS techniques in underwater environments. For instance, [8] demonstrated the feasibility of using backscattered signals for long-distance passive communications with a Van Atta array. By switching the load of each reflector between short and open circuit states, the polarization of the reflected wave is toggled. Consequently, the strength of the received signal, which is the superposition of emitted and reflected waves, exhibits variations. Sea experiments indicate that the Van Atta backscatter array can achieve a round-trip communication range exceeding 300 meters with a bit error rate (BER) of $10^{-3}$.
@@
In [21], the enhancement of communication performance using UA-RIS was examined. In this design, each reflection unit incorporates a piezoelectric coil positioned between two metal plates, with the resonant frequency adjusted by the reverse voltage applied to a varactor. Meanwhile, [27] explored the design of UA-RIS capable of responding to a wide range of frequencies. This work addressed element dispersion within each reflector and array dispersion throughout the acoustic RIS. By counteracting these dispersions, the directivity of the reflected waves becomes frequency-independent, allowing wideband signals to achieve uniform beam patterns across various frequency components. The efficacy of this architecture was validated through Bellhop and COMSOL simulators.
@@ III Motivation of Multilayered UA-RIS
This section outlines the motivation behind adopting a multilayered UA-RIS. We begin by discussing the importance of simultaneously manipulating both the amplitude and phase of reflected waves in UA-RIS applications. Afterward, the inherent challenges in achieving flexible reflection control are clarified. Finally, we briefly review our previous solution for implementing such control and highlight its limitations.
@@ III-A Necessity of Flexible Reflection Control
The underwater acoustic channel is a crucial resource, vital not only for manmade facilities but also for marine animals that rely on acoustic signals for hunting, communication, and navigation [14]. However, the absorption attenuation of sound signals increases significantly with frequency, limiting the available bandwidth for mid-range and long-range acoustic applications to just tens of kilohertz. As the number of marine facilities grows, the need to efficiently utilize the acoustic channel without adversely impacting the surrounding marine life becomes a critical challenge for future acoustic systems.
To meet the above objective, multiple UA-RIS facilities can be deployed on the seafloor at predetermined locations. As illustrated in Fig. 1, each facility is connected to a wirewalker [17], a fully passive device that harvests energy from ocean-wave motion. With appropriate system design, UA-RIS units can steer the source signal toward the intended receiver and coherently combine the reflected acoustic waves, thereby significantly enhancing the signal quality. In addition, every UA-RIS can naturally work as a passive sonar system. As a result, the relative directions of the sender-reflector and receiver-reflector paths can be estimated using angle-of-arrival (AoA) measurements.
Figure 1: Scenario of UA-RIS assisted underwater acoustic systems for eco-friendly communications.
The goal of UA-RIS is to introduce a controllable, phase-adjustable propagation component whose amplitude and delay are engineered to (a) combine constructively with the existing arrivals at the receiver, and (b) provide an additional high-SNR path that the receiver can either resolve or coherently combine to enhance overall communication performance. This capability will allow the source level of acoustic transmissions to be reduced without degrading communication performance, ultimately supporting more environmentally sustainable underwater communication.
To achieve optimal performance, RIS-assisted underwater acoustic communication systems must be designed to not only maximize signal strength in the main lobe but also suppress interference arising from side lobes. The typical 1-bit and 2-bit coding schemes used in RF-RIS may fall short, as the phase resolution of RIS elements is a significant determinant of system performance. According to array processing principles [26], an RIS with arbitrary reflection coefficients can increase the directional gain of the reflected wave by approximately 23.46% ($\sim$ 0.9 dB) over a 2-bit coding RIS and diminish side lobe interference by around 50% ($\sim$ 3 dB).
@@
III-B Challenges of Flexible Reflection Control in UA-RIS
In RF-RIS, the flexible phase shift is achieved by adjusting the reverse bias voltage applied to the varactor’s p-n junction [18, 1]. However, this approach is impractical for UA-RIS due to the acoustic signals operating at frequencies six orders of magnitude lower than RF communications. Specifically, the reflection coefficient, denoted as $\Gamma$, in a transmission line is defined as:
@@ III-B Challenges of Flexible Reflection Control in UA-RIS
For instance, adjusting the phase of  $\Gamma$ from -45 to -90 requires changing the load impedance from j-j121 $\Omega$ to j-j50 $\Omega$. At the 2.4 GHz RF frequency, the capacitance of the load capacitor needs to vary from 0.55 pF to 1.33 pF. This range falls within the capability of many commercial varactor diodes [16, 31]. In contrast, if the frequency is reduced to 25 kHz for long-distance acoustic communication, the required capacitance of the load capacitor must change from 52 nF to 128 nF to achieve the same phase adjustment. This range significantly exceeds the capabilities of varactor diodes, which typically have capacitance values below 1 nF.
III-C Previous Solution of Flexible Reflection Control and its Limitations
@@
In our previous work [13], we introduced a synthetic reflection-based RIS to synthesize acoustic waves with customizable amplitude and phase. This approach utilized two reflectors to generate in-phase and quadrature components. In contrast, the new multilayer structure proposed in this paper requires only one reflector. We will next provide a detailed comparison of the two methods.
Figure 2: Synthetic reflection implemented with separated reflectors.
@@
Fig. 2 illustrates the synthetic reflection scheme from [13], where two adjacent reflectors sharing the same Y-coordinate are paired. For example, reflector 5 paired with reflector 6, collectively act as a single unit on reflection plane 1. This is possible because their projected positions along the Y-axis are aligned, ensuring that positional differences do not introduce additional phase shifts on that plane. During operation, these paired reflectors produce orthogonal reflection coefficients, allowing the synthesis of reflected waves with arbitrary phases on plane 1 through appropriate adjustments of the amplitude ratio between these orthogonal components.
@@
Using this method, two groups of reflectors indicated by purple and blue in Fig. 2 are formed. Each group consists of four reflector pairs, forming four collective reflectors along the Y-axis. These collective reflectors can generate flexible amplitude and phase of the reflection coefficients. By varying the phase differences among neighboring collective reflectors, the angle of reflection plane 1 can be effectively altered.
@@
In an ideal scenario, the reflected wave would be superposed along a single line oriented toward the target, instead of spreading over a plane. As illustrated in Fig. 2, this is achieved by configuring two sets of linear arrays along the X and Y axis to form two reflection planes, with their intersection defining the array’s directivity.
@@
However, in the UA-RIS proposed in our previous research [13], all reflection coefficients are determined solely by the angle of the reflection plane 1 and the incident wave, leaving no additional DoF to establish a second reflection plane, leaving no additional degrees of freedom to form a second reflection plane. This design limitation prevents pairing reflectors along both the X and Y axes, restricting synthetic reflection to one-dimensional gain without enhancing directivity in the orthogonal dimension. For an $N\!\times\!N$ array, this reduces the performance to that of $N/2$ independent linear arrays, each with NN acoustic reflectors.
@@
To improve UA-RIS performance, a new structure is necessary to flexibly control the phase and the amplitude of reflected waves. The criteria for this structure are: (a) The design should achieve optimal or near-optimal main lobe gain and side lobe suppression, approaching the performance of an ideal two-dimensional RIS. (b) It should readily adapt to varying environmental conditions, such as variation in intensity and angle of the incident wave, temperature changes, and deployment depths, thereby expanding the application range of UA-RIS. (c) The new structure should be economically feasible, enabling large-scale deployment.